Because L^{A}TEX will almost always convert a sequence of spaces into a single space, it can be rather difficult to lay out tables. See what happens in this example

     \begin{flushleft}
     Income  Expenditure Result   \\
     20s 0d  19s 11d     happiness \\
     20s 0d  20s 1d      misery  \\
     \end{flushleft}

Income Expenditure Result 20s 0d 19s 11d happiness 20s 0d 20s 1d misery

The tabbing environment overcomes this problem. Within it you set tabstops and tab to them much like you do on a typewriter. Tabstops are set with the \= command, and the \> command moves to the next stop. The \\ command is used to separate each line. A line that ends \kill produces no output, and can be used to set tabstops:

     \begin{tabbing}
     Income \=Expenditure \=    \kill
     Income \>Expenditure \>Result \\
     20s 0d \>19s 11d \>Happiness   \\
     20s 0d \>20s 1d  \>Misery    \\
     \end{tabbing}

Income Expenditure Result 20s 0d 19s 11d Happiness 20s 0d 20s 1d Misery

Unlike a typewriter’s tab key, the \> command always moves to the next tabstop in sequence, even if this means moving to the left. This can cause text to be overwritten if the gap between two tabstops is too small.